The amount of flux in a round magnetic rod measured at 0.013 weber.If the diameter of the material is 12 cm, calculate the flux density؟
1
Expert's answer
2020-10-12T11:43:59-0400

The flux density is


"\phi=\frac{\Phi}{A}=\frac{4\Phi}{\pi d^2},\\\space\\
\phi=\frac{4\cdot0.013}{\pi\cdot0.12^2}=1.15\text{ Wb/m}^2."